# Minimmit: Fast Finality with Even Faster Blocks

### v1

@@ 4. The formal specification
We initially give a specification aimed at simplicity. In the full version of the paper, we will describe optimisations, such as the use of threshold signatures to reduce communication complexity. In what follows, we suppose that all messages are signed by the sender, and that, when a correct processor sends a message to ‘all processors’, it regards that message as immediately received by itself. The pseudocode uses a number of message types, local variables, functions and procedures, detailed below.
@@
M-notarizations. An M-notarization for the block $b$ is a set of $2f+1$ votes for $b$, each signed by a different processor. (By an M-notarization, we mean an M-notarization for some block.)
L-notarizations. An L-notarization for the block $b$ is a set of nfn-f votes for $b$, each signed by a different processor.
@@
The local variable $\mathtt{S}$. This variable is maintained locally by each processor $p_{i}$ and stores all messages received. It is considered to be automatically updated, i.e., we do not give explicit instructions in the pseudocode updating $\mathtt{S}$. We say a set of messages $M^{\prime}$ is contained in $\mathtt{S}$ if $M^{\prime}\subseteq\mathtt{S}$. We also regard $\mathtt{S}$ as containing a block $b$ whenever $\mathtt{S}$ contains any message (tuple) with $b$ as one of its entries. Initially, $\mathtt{S}$ contains only $b_{\text{gen}}$ and an L-notarization (and an M-notarization) for $b_{\text{gen}}$.
@@
    An M-notarization for some $b^{\prime}$ with $H(b^{\prime})=h$, and with $b^{\prime}.\text{view}=v^{\prime}$ (say), and;
@@
New nullifications and notarizations. Processors will be required to forward on all newly received nullifications and notarizations. To make this precise, we must specify what is to count as a ‘new’ nullification/notarization. At timeslot $t$, $p_{i}$ regards a nullification $N\subseteq\mathtt{S}$ for some view $v$ (not necessarily equal to $\mathtt{v}$) as new if:
@@
At timeslot $t$, $p_{i}$ regards an $M$-notarization $Q\subseteq\mathtt{S}$ for block $b$ as new if:
  • $\mathtt{S}$ did not contain an M-notarization for $b$ at any smaller timeslot, and;
  • $Q$ is lexicographically least amongst M-notarizations for $b$ contained in $\mathtt{S}$.
Similarly, $p_{i}$ regards an L-notarization $Q\subseteq\mathtt{S}$ for block $b$ as new if the two bullets above hold with ‘M-notarization’ replaced by ‘L-notarization’.
@@
M-notarizaton for $b$ A set of $2f+1$ votes for $b$, each signed by a different processor
L–notarizaton for $b$ A set of nfn-f votes for $b$, each signed by a different processor
@@
$\mathtt{notarized}$ Initially set to $\bot$, records block voted for in present view
@@
Initially contains only $b_{\text{gen}}$ and M/L-notarizations for $b_{\text{gen}}$
@@
3:   Send new M- and L-notarizations in $\mathtt{S}$ to all processors;
@@
5:   If $p_{i}=\mathtt{lead}(\mathtt{v})$:
@@
9:    If $\mathtt{notarized}=\bot$ and $\mathtt{nullified}=$false:
10:    Set $\mathtt{notarized}:=b$ and send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
@@
12:   If $\mathtt{T}=2\Delta$, $\mathtt{nullified}=$false and $\mathtt{notarized}=\bot$:
@@
16:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$false, $\mathtt{notarized}:=\bot$; $\triangleright$ Go to next view
17:   If $\mathtt{S}$ contains an M-notarization for some $b$ with $b.\text{view}=\mathtt{v}$:
18:    If $\mathtt{notarized}=\bot$ and $\mathtt{nullified}=$false, send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
19:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$false, $\mathtt{notarized}:=\bot$; $\triangleright$ Go to next view
@@
21:   If $\mathtt{nullified}=$false, $\mathtt{notarized}\neq\bot$ and $\mathtt{S}$ contains $\geq 2f+1$ messages, each signed by a
@@
24:   (ii) Of the form $(\text{vote},b)$ for some $b$ s.t. $b.\text{view}=\mathtt{v}$ and $\mathtt{notarized}\neq b$:
@@
28:   If $\mathtt{S}$ contains a new L-notarization for any block $b$:
@@ 5.1. Consistency
We say block $b$ receives an M-notarization if $b=b_{\text{gen}}$ or at least $2f+1$ processors send votes for $b$. Similarly, we say $b$ receives an L-notarization if $b=b_{\text{gen}}$ or at least nfn-f processors send votes for $b$. View $v$ receives a nullification if at least $2f+1$ processors send nullify(v)(v) messages.
@@
Recall that $\bot$ is a default value, different than any block. Each correct processor’s local value $\mathtt{notarized}$ is initially set to $\bot$, and is also set to $\bot$ upon entering any view (lines 16 and 19). A correct processor $p_{i}$ will only vote for a block if $\mathtt{notarized}=\bot$ (lines 9 and 18). The claim of the lemma holds because, upon voting for any block $b$, $p_{i}$ either sets $\mathtt{notarized}:=b$ (line 10) and then does not redefine this value until entering the next view, or else immediately enters the next view (lines 18 and 19). ∎
@@
If $b$ receives an L-notarization, then no block $b^{\prime}\neq b$ with $b^{\prime}.\text{view}=b.\text{view}$ receives an M-notarization.
@@
Given Lemma 5.1, this now follows as in Section 3. Towards a contradiction, suppose that $b$ receives an L-notarization and that $b^{\prime}\neq b$ with $b^{\prime}.\text{view}=b.\text{view}$ receives an M-notarization. Let $P$ be the set of processors that contribute to the L-notarization for $b$, and let $P^{\prime}$ be the set of processors that vote for $b^{\prime}$. Then $|P\cap P^{\prime}|\geq(n-f)+(2f+1)-n=f+1$. So, $P\cap P^{\prime}$ contains a correct processor, which contradicts Lemma 5.1. ∎
@@
If $b$ receives an L-notarization and $v=b.\text{view}$, then view $v$ does not receive a nullification.
@@
Towards a contradiction, suppose $b$ receives an L-notarization, $v=b.\text{view}$, and view $v$ receives a nullification. Let $P$ be the correct processors that vote for $b$, let $P^{\prime}=\Pi\setminus P$, and note that $|P^{\prime}|\leq 2f$. Since view $v$ receives a nullification, it follows that some processor in $P$ must send a nullify(v)(v) message. So, let $t$ be the first timeslot at which some processor $p_{i}\in P$ sends such a message. Since $p_{i}$ cannot send a nullify(v)(v) message upon timeout (lines 12-13), $p_{i}$ must send the nullify(v)(v) message at $t$ because the conditions of lines 21-24 hold for $p_{i}$ at $t$, i.e., $p_{i}$ must have received $\geq 2f+1$ messages, each signed by a different processor, and each of the form:
@@
Towards a contradiction, suppose that two inconsistent blocks, $b$ and $b^{\prime}$ say, both receive L-notarizations. Without loss of generality suppose $b.\text{view}\leq b^{\prime}.\text{view}$. Set $b_{1}:=b$ and $v_{1}:=b_{1}.\text{view}$. Then there is a least $v_{2}\geq v_{1}$ such that some block $b_{2}$ satisfies:
@@
    $b_{2}$ receives an M-notarization.
From Lemma 5.2, it follows that $v_{2}>v_{1}$. According to clause (ii) from the definition of when $\mathtt{S}$ contains a valid proposal for view $v_{2}$, correct processors will not vote for $b_{2}$ in line 10 until they receive an M-notarization for its parent, $b_{0}$ say. Correct processors will not vote for $b_{2}$ in line 18 until $b_{2}$ has already received an M-notarization, meaning that at least $f+1$ correct processors must first vote for $b_{2}$ via line 10, and $b_{0}$ must receive an M-notarization. By our choice of $v_{2}$, it follows that $b_{0}.\text{view}<v_{1}$. This gives a contradiction, because, by clause (iii) from the definition of a valid proposal for view $v_{2}$, correct processors would not vote for $b_{2}$ in line 10 without receiving a nullification for view $v_{1}$. By Lemma 5.3, such a nullification cannot exist. So, block $b_{2}$ cannot receive an M-notarization (and no correct processor votes for $b_{2}$ via either line 10 or 18). ∎
@@ 5.2. Liveness
Towards a contradiction, suppose that some correct processor $p_{i}$ enters view $v$, but never enters view $v+1$. Note that correct processors only leave any view $v^{\prime}$ upon receiving either a nullification for the view, or else an M-notarization for some view $v^{\prime}$ block. Since correct processors forward new nullifications and notarizations upon receiving them (lines 2 and 3), the fact that $p_{i}$ enters view $v$ but does not leave it means that:
@@
Each correct processor eventually receives, from at least nfn-f processors, either a vote for some view $v$ block, or a nullify(v)(v) message. If any correct processor receives an M-notarization for a view $v$ block, then we reach an immediate contradiction. So, suppose otherwise. If $p_{j}$ is a correct processor that votes for a view $v$ block $b$, it follows that $p_{j}$ receives messages from at least $(n-f)-(2f)=n-3f\geq 2f+1$ processors, each of which is either:
@@
If $p_{i}=\mathtt{lead}(v)$ is correct, and if the first correct processor to enter view $v$ does so after GST, then $p_{i}$ sends a block to all processors and that block receives an L-notarization.
@@
Suppose $p_{i}=\mathtt{lead}(v)$ is correct and that the first correct processor $p_{j}$ to enter view $v$ does so at timeslot $t\geq\text{GST}$. If $v>1$, processor $p_{j}$ enters view $v$ upon receiving either a nullification for view v1v-1, or else an M-notarization for some view v1v-1 block. Since $p_{j}$ forwards on all new notarizations and nullifications that it receives (lines 2 and 3), it follows that all correct processors enter view $v$ by $t+\Delta$ (note that this also holds if $v=1$). Processor $p_{i}$ therefore sends a new block $b$ to all processors by $t+\Delta$, which is received by all processors by $t+2\Delta$. Let $b^{\prime}$ be the parent of $b$ and suppose $b^{\prime}.\text{view}=v^{\prime}$. Then $p_{i}$ receives an M-notarization for $b^{\prime}$ by $t+\Delta$. Since $p_{i}$ forwards on all new notarizations that it receives (line 3), all correct processors receive an M-notarization for $b^{\prime}$ by $t+2\Delta$. Since $p_{i}$ has entered view $v$, it must also have received nullifications for all views in the open interval $(v^{\prime},v)$ by $t+\Delta$, and all correct processors receive these by $t+2\Delta$. All correct processors therefore vote for $b$ (by either line 10 or 18) [^8] The point of line 18 is to ensure this part of the proof goes through. Without it, there is the possibility that correct processors move to the next view upon seeing an M-notarization, before they are able to vote via line 10, thereby failing to guarantee an L-notarization. before any correct processor sends a nullify(v)(v) message. The block $b$ therefore receives an L-notarization, as claimed. ∎

### v2

@@ 4. The formal specification
We initially give a specification aimed at simplicity. In Section 6, we will describe optimisations, such as the use of threshold signatures to reduce communication complexity. In what follows, we suppose that all messages are signed by the sender, and that, when a correct processor sends a message to ‘all processors’, it regards that message as immediately received by itself. For the sake of simplicity, we also initially assume (without explicit mention in the pseudocode) that correct processors automatically send new transactions to all others upon first receiving them - we will revisit this assumption in Section 6. The pseudocode uses a number of message types, local variables, functions and procedures, detailed below.
@@
M-notarisations. An M-notarisation for the block $b$ is a set of $2f+1$ votes for $b$, each signed by a different processor. (By an M-notarisation, we mean an M-notarisation for some block.)
L-notarisations. An L-notarisation for the block $b$ is a set of nfn-f votes for $b$, each signed by a different processor. We note that, since an L-notarisation is a larger set of votes than an M-notarisation, if $p_{i}$ has received an L-notarisation for $b$, then it has necessarily received an M-notarisation for $b$.
@@
The local variable $\mathtt{S}$. This variable is maintained locally by each processor $p_{i}$ and stores all messages received. It is considered to be automatically updated, i.e., we do not give explicit instructions in the pseudocode updating $\mathtt{S}$. We say a set of messages $M^{\prime}$ is contained in $\mathtt{S}$ if $M^{\prime}\subseteq\mathtt{S}$. We also regard $\mathtt{S}$ as containing a block $b$ whenever $\mathtt{S}$ contains any message (tuple) with $b$ as one of its entries. Initially, $\mathtt{S}$ contains only $b_{\text{gen}}$ and an L-notarisation (and an M-notarisation) for $b_{\text{gen}}$.
@@
    An M-notarisation for some $b^{\prime}$ with $H(b^{\prime})=h$, and with $b^{\prime}.\text{view}=v^{\prime}$ (say), and;
@@
New nullifications and notarisations. Processors will be required to forward on all newly received nullifications and M-notarisations. To make this precise, we must specify what is to count as a ‘new’ nullification/notarisation. At timeslot $t$, $p_{i}$ regards a nullification $N\subseteq\mathtt{S}$ for some view $v$ (not necessarily equal to $\mathtt{v}$) as new if:
@@
At timeslot $t$, $p_{i}$ regards an $M$-notarisation $Q\subseteq\mathtt{S}$ for block $b$ as new [^10] Since it is not necessary for liveness, our pseudocode does not require processors to forward L-notarisations. One could also require processors to forward L-notarisations, and the proofs of Section 5 would go through unchanged. if:
  • $\mathtt{S}$ did not contain an M-notarisation for $b$ at any smaller timeslot, and;
  • $Q$ is lexicographically least amongst M-notarisations for $b$ contained in $\mathtt{S}$.
@@
M-notarisation for $b$ A set of $2f+1$ votes for $b$, each signed by a different processor
L-notarisation for $b$ A set of nfn-f votes for $b$, each signed by a different processor
@@
$\mathtt{proposed}$ Initially false, specifies whether already proposed a block for view $v$
$\mathtt{notarised}$ Initially set to $\bot$, records block voted for in present view
@@
Initially contains only $b_{\text{gen}}$ and M/L-notarisations for $b_{\text{gen}}$
@@
3:   Send new M-notarisations in $\mathtt{S}$ to all processors;
@@
5:   If $p_{i}=\mathtt{lead}(\mathtt{v})$ and $\mathtt{proposed}=$ false:
@@
9:    If $\mathtt{notarised}=\bot$ and $\mathtt{nullified}=$ false:
10:    Set $\mathtt{notarised}:=b$ and send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
@@
12:   If $\mathtt{T}=2\Delta$, $\mathtt{nullified}=$ false and $\mathtt{notarised}=\bot$:
@@
16:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$ false, $\mathtt{proposed}=$ false, $\mathtt{notarised}:=\bot$; $\triangleright$ Go to next view
17:   If $\mathtt{S}$ contains an M-notarisation for some $b$ with $b.\text{view}=\mathtt{v}$:
18:    If $\mathtt{notarised}=\bot$ and $\mathtt{nullified}=$ false, send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
19:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$ false, $\mathtt{proposed}=$ false, $\mathtt{notarised}:=\bot$; $\triangleright$ Go to next view
@@
21:   If $\mathtt{nullified}=$ false, $\mathtt{notarised}\neq\bot$ and $\mathtt{S}$ contains $\geq 2f+1$ messages, each signed by a
@@
24:   (ii) Of the form $(\text{vote},b)$ for some $b$ s.t. $b.\text{view}=\mathtt{v}$ and $\mathtt{notarised}\neq b$:
@@
28:   If $\mathtt{S}$ contains a new L-notarisation for any block $b$:
@@ 5.1. Consistency
We say block $b$ receives an M-notarisation if $b=b_{\text{gen}}$ or at least $2f+1$ processors send votes for $b$. Similarly, we say $b$ receives an L-notarisation if $b=b_{\text{gen}}$ or at least nfn-f processors send votes for $b$. View $v$ receives a nullification if at least $2f+1$ processors send nullify(v)(v) messages.
@@
Recall that $\bot$ is a default value, different than any block. Each correct processor’s local value $\mathtt{notarised}$ is initially set to $\bot$, and is also set to $\bot$ upon entering any view (lines 16 and 19). A correct processor $p_{i}$ will only vote for a block if $\mathtt{notarised}=\bot$ (lines 9 and 18). The claim of the lemma holds because, upon voting for any block $b$, $p_{i}$ either sets $\mathtt{notarised}:=b$ (line 10) and then does not redefine this value until entering the next view, or else immediately enters the next view (lines 18 and 19). ∎
@@
If $b$ receives an L-notarisation, then no block $b^{\prime}\neq b$ with $b^{\prime}.\text{view}=b.\text{view}$ receives an M-notarisation.
@@
Given Lemma 5.1, this now follows as in Section 3. Towards a contradiction, suppose that $b$ receives an L-notarisation and that $b^{\prime}\neq b$ with $b^{\prime}.\text{view}=b.\text{view}$ receives an M-notarisation. Let $P$ be the set of processors that contribute to the L-notarisation for $b$, and let $P^{\prime}$ be the set of processors that vote for $b^{\prime}$. Then $|P\cap P^{\prime}|\geq(n-f)+(2f+1)-n=f+1$. So, $P\cap P^{\prime}$ contains a correct processor, which contradicts Lemma 5.1. ∎
@@
If $b$ receives an L-notarisation and $v=b.\text{view}$, then view $v$ does not receive a nullification.
@@
Towards a contradiction, suppose $b$ receives an L-notarisation, $v=b.\text{view}$, and view $v$ receives a nullification. Let $P$ be the correct processors that vote for $b$, let $P^{\prime}=\Pi\setminus P$, and note that $|P^{\prime}|\leq 2f$. Since view $v$ receives a nullification, it follows that some processor in $P$ must send a nullify(v)(v) message. So, let $t$ be the first timeslot at which some processor $p_{i}\in P$ sends such a message. Since $p_{i}$ cannot send a nullify(v)(v) message upon timeout (lines 12-13), $p_{i}$ must send the nullify(v)(v) message at $t$ because the conditions of lines 21-24 hold for $p_{i}$ at $t$, i.e., $p_{i}$ must have received $\geq 2f+1$ messages, each signed by a different processor, and each of the form:
@@
Towards a contradiction, suppose that two inconsistent blocks, $b$ and $b^{\prime}$ say, both receive L-notarisations. Without loss of generality suppose $b.\text{view}\leq b^{\prime}.\text{view}$. Set $b_{1}:=b$ and $v_{1}:=b_{1}.\text{view}$. Then there is a least $v_{2}\geq v_{1}$ such that some block $b_{2}$ satisfies:
@@
    $b_{2}$ receives an M-notarisation.
From Lemma 5.2, it follows that $v_{2}>v_{1}$. According to clause (ii) from the definition of when $\mathtt{S}$ contains a valid proposal for view $v_{2}$, correct processors will not vote for $b_{2}$ in line 10 until they receive an M-notarisation for its parent, $b_{0}$ say. Correct processors will not vote for $b_{2}$ in line 18 until $b_{2}$ has already received an M-notarisation, meaning that at least $f+1$ correct processors must first vote for $b_{2}$ via line 10, and $b_{0}$ must receive an M-notarisation. By our choice of $v_{2}$, it follows that $b_{0}.\text{view}<v_{1}$. This gives a contradiction, because, by clause (iii) from the definition of a valid proposal for view $v_{2}$, correct processors would not vote for $b_{2}$ in line 10 without receiving a nullification for view $v_{1}$. By Lemma 5.3, such a nullification cannot exist. So, block $b_{2}$ cannot receive an M-notarisation (and no correct processor votes for $b_{2}$ via either line 10 or 18). ∎
@@ 5.2. Liveness
Towards a contradiction, suppose that some correct processor $p_{i}$ enters view $v$, but never enters view $v+1$. Note that correct processors only leave any view $v^{\prime}$ upon receiving either a nullification for the view, or else an M-notarisation for some view $v^{\prime}$ block. Since correct processors forward new nullifications and notarisations upon receiving them (lines 2 and 3), the fact that $p_{i}$ enters view $v$ but does not leave it means that:
@@
Each correct processor eventually receives, from at least nfn-f processors, either a vote for some view $v$ block, or a nullify(v)(v) message. If any correct processor receives an M-notarisation for a view $v$ block, then we reach an immediate contradiction. So, suppose otherwise. If $p_{j}$ is a correct processor that votes for a view $v$ block $b$, it follows that $p_{j}$ receives messages from at least $(n-f)-(2f)=n-3f\geq 2f+1$ processors, each of which is either:
@@
If $p_{i}=\mathtt{lead}(v)$ is correct, and if the first correct processor to enter view $v$ does so after GST, then $p_{i}$ sends a block to all processors and that block receives an L-notarisation.
@@
Suppose $p_{i}=\mathtt{lead}(v)$ is correct and that the first correct processor $p_{j}$ to enter view $v$ does so at timeslot $t\geq\text{GST}$. If $v>1$, processor $p_{j}$ enters view $v$ upon receiving either a nullification for view v1v-1, or else an M-notarisation for some view v1v-1 block. Since $p_{j}$ forwards on all new notarisations and nullifications that it receives (lines 2 and 3), it follows that all correct processors enter view $v$ by $t+\Delta$ (note that this also holds if $v=1$). Processor $p_{i}$ therefore sends a new block $b$ to all processors by $t+\Delta$, which is received by all processors by $t+2\Delta$. Let $b^{\prime}$ be the parent of $b$ and suppose $b^{\prime}.\text{view}=v^{\prime}$. Then $p_{i}$ receives an M-notarisation for $b^{\prime}$ by $t+\Delta$. Since $p_{i}$ forwards on all new notarisations that it receives (line 3), all correct processors receive an M-notarisation for $b^{\prime}$ by $t+2\Delta$. Since $p_{i}$ has entered view $v$, it must also have received nullifications for all views in the open interval $(v^{\prime},v)$ by $t+\Delta$, and all correct processors receive these by $t+2\Delta$. All correct processors therefore vote for $b$ (by either line 10 or 18) [^11] The point of line 18 is to ensure this part of the proof goes through. As noted in Section 3, without it, there is the possibility that correct processors move to the next view upon seeing an M-notarisation, before they are able to vote via line 10, thereby failing to guarantee an L-notarisation. before any correct processor sends a nullify(v)(v) message. The block $b$ therefore receives an L-notarisation, as claimed. ∎